OpenStudy (thickgirl_angel):

The variable Z is directly proportional to X. When X is 12, Z has the value 228. What is the value of Z when X = 18

OpenStudy (zzr0ck3r):

\(x=zk\) \(12=228*k \implies k=\frac{12}{228}\) So \(x=z*\frac{12}{228}\) Plug in \(x=18\) to find the corresponding \(z\)

OpenStudy (radar):

@zz0ck3r The stated relationship between the variable Z and the variable X can be expressed as: Z = kX. This seems to be clearer than X = kZ. Simply because the problem does not state that X is directly proportional to Z. Although obviously it would have to be.